# HyperFusion: A Hypernetwork Approach to Multimodal Integration of Tabular and Medical Imaging Data for Predictive Modeling

### v1

@@ Abstract
The integration of diverse clinical modalities such as medical imaging and the tabular data obtained by the patients’ Electronic Health Records (EHRs) is a crucial aspect of modern healthcare. The integrative analysis of multiple sources can provide a comprehensive understanding of a patient’s condition and can enhance diagnoses and treatment decisions. Deep Neural Networks (DNNs) consistently showcase outstanding performance in a wide range of multimodal tasks in the medical domain. However, the complex endeavor of effectively merging medical imaging with clinical, demographic and genetic information represented as numerical tabular data remains a highly active and ongoing research pursuit.
We present a novel framework based on hypernetworks to fuse clinical imaging and tabular data by conditioning the image processing on the EHR’s values and measurements. This approach aims to leverage the complementary information present in these modalities to enhance the accuracy of various medical applications. We demonstrate the strength and the generality of our method on two different brain Magnetic Resonance Imaging (MRI) analysis tasks, namely, brain age prediction conditioned by subject’s sex, and multiclass Alzheimer’s Disease (AD) classification conditioned by tabular data. We show that our framework outperforms both single-modality models and state-of-the-art MRI-tabular data fusion methods. The code, enclosed to this manuscript will be made publicly available.
@@ 1 Introduction
In medical decision-making, clinicians depend on comprehensive patient information to make precise diagnoses and well-informed decisions. The refinement and amalgamation of all available data are crucial for personalized healthcare. This is particularly relevant in image-based diagnoses, where the inclusion of clinical and demographic patient information from the Electronic Health Record (EHR) significantly improves diagnostic accuracy.
Deep Neural Networks (DNNs) consistently achieve state-of-the-art results across a broad spectrum of medical tasks. However, they still cannot match human experts’ abilities to fuse medical imaging and EHR data, typically presented in tabular format. The main challenge arises from the inherent disparities between these two types of data. Images are high-dimensional, continuous, and spatial, while tabular data typically comprise low-dimensional features with diverse types, scales, and ranges. Nevertheless, employing DNNs for this integration offers significant advantages, as they can process extensive volumes of data and extract intricate structural relationships. The proposed research aims to leverage DNNs’ capabilities while addressing the challenges of data fusion through hyper-networks.
When considering multi-modal fusion of imaging and non-imaging data, the integration of images and textual data comes to mind. This well-studied problem has attracted significant attention following its outstanding success (Perez et al., 2018; Anderson et al., 2018; Radford et al., 2021). In the clinical domain, numerous methods have been proposed to combine text and medical imaging (Wang et al., 2022). However, while texts significantly differ from images in terms of structure and dimension, these two sources of data often have explicit semantic relations, as seen in image captions (Vinyals et al., 2015; Xu et al., 2015; Wang et al., 2016), generation of images from text (Zhang et al., 2023, 2018), visual question answering (Perez et al., 2018; Anderson et al., 2018), or radiology interpretation (Dawidowicz et al., 2023; Wang et al., 2022).
Dealing with tabular data, certain measures (typically related to size or morphology) may directly correspond to the scanned anatomy. Recent work by Hager et al. has utilized such correspondences to establish associations between different modalities (Hager et al., 2023). However, numerous other clinical and demographic attributes, including measures obtained in medical laboratory tests, the medical history of patients or their biological relatives, and even years of education, provide complementary information, independent of the imaging data. For example, a chest X-ray may require additional clinical measures (e.g., body temperature) for an accurate diagnosis of pneumonia (Huang et al., 2020). Our goal is to develop a deep learning model capable of enhancing data utilization by extracting implicit, possibly complex correlations between different, modality-specific characteristics.
A straightforward and intuitive approach to fuse image and tabular data involves using separate networks for each data type and making a collective decision based on their outcomes (Prabhu et al., 2022). The processing of each data type typically involves a Multi-Layer Perceptron (MLP) for tabular data and a Convolutional Neural Network (CNN) for imaging. Another common approach entails embedding each modality into a feature vector followed by vector concatenation (Esmaeilzadeh et al., 2018; El-Sappagh et al., 2020; Venugopalan et al., 2021). Other methods, such as those presented in (Wolf et al., 2022), leverage tabular data to condition image processing through affine transformations in the convolutional layers. In all of these techniques, the interaction between different data types is confined to specific stages of the processing pipeline and specific transformations, potentially limiting the network’s ability to fully exploit the advantages of data fusion.
@@
We present a conceptually innovative approach to the challenge of fusing imaging and tabular data. Specifically, we designate tabular measurements and values as priors influencing the outcomes of the image analysis network. This can be likened to conditioning the results of the image analysis network on the tabular data, comparable to adjusting the auditory output of a radio transmitter by manipulating frequency (selecting stations) and volume. The central concept is implemented through hypernetworks. Originally introduced by (Ha et al., 2016), the hypernetwork framework comprises a primary network assigned to a specific task and a hypernetwork responsible for generating weights and biases for specific layers in the primary network. This arrangement enables the framework to condition the processing of one modality in the primary network based on information from another modality processed within the hypernetwork. The entire fusion process is akin to finely adjusting the focus of image processing based on attributes present in the tabular data.
We demonstrate the versatility and efficacy of the proposed hypernetwork framework, named HyperFusion, through two distinct brain MRI analysis tasks: brain age prediction conditioned by the subject’s sex and classification of subjects into Alzheimer’s disease (AD), Mild Cognitive Impairment (MCI), and Cognitively Normal (CN) groups conditioned by their tabular data, which includes clinical measurements, as well as demographic and genetic information.
@@
The rest of this paper is structured as follows: In Section 2, we present a literature review on imaging and tabular data fusion. We also delve into the concept of hypernetworks and their applications. Furthermore, we discuss studies related to the brain age prediction and AD classification tasks. Section 3 presents our hypernetwork framework along with details about the architectures designed for each task and key implementation aspects. In Section 4, we describe our experimental setup, including the datasets used and preprocessing steps, as well as training and evaluation procedures. The results of our experiments are presented and discussed in this section as well. We conclude in Section 5.
@@ 2.1 Fusion methods
The emergence of public multi-modal medical imaging datasets and benchmarks have facilitated the development of methods for integrating structural images of different contrasts for enhanced clinical analysis and, in particular, pathology segmentation (Menze et al., 2014; Carass et al., 2017). Recent efforts concentrate on accommodating missing imaging (mainly MRI) modalities during training time and inference, which appears to be a main bottleneck (Yang et al., 2023). Another line of works focuses on the synergy between functional, structural and diffusion images which require additional processing since the voxel values of each modality convey different properties (Sui et al., 2023). However, the incorporation of modalities with no spatial structure, such as tabular data, poses further challenges.
Fusion strategies for combining medical imaging and tabular data can be categorized into three distinct types, as outlined in (Huang et al., 2020): early fusion, where original or extracted features are concatenated at the input level, e.g., (Chieregato et al., 2022); joint fusion (or intermediate fusion), where the feature extraction phase is learned as part of the fusion model; and late fusion, where predictions or pre-trained high level features are combined at the decision level, as in (Pandeya and Lee, 2021; Prabhu et al., 2022).
@@
A straightforward joint fusion approach for integrating imaging and tabular data involves concatenating their latent feature vectors, as exemplified in studies such as (Esmaeilzadeh et al., 2018; El-Sappagh et al., 2020; Venugopalan et al., 2021). In this approach, tabular data or processed tabular features are merged with high-level image features through a fully connected layer or a small MLP, enabling gradient propagation to both tabular and image processing networks. However, methods based on late concatenation may limit the interaction between tabular data and imaging to high-level descriptors, neglecting fusion in convolutional layers where crucial spatial image context is preserved.
A more advanced joint fusion approach operates within the intermediate layers of a CNN by applying an affine transformation to the image features. The shift and scale parameters are generated by a network that processes another data modality. In (Perez et al., 2018), a Feature-wise Linear Modulation (FiLM) was introduced for conditioning image processing on text, applied to a visual question-answering task. Building upon FiLM, (Wolf et al., 2022) introduced the Dynamic Affine Feature Map Transform (DAFT) to fuse brain MRIs and tabular data. We compared our framework to both DAFT and FiLM. It is worth noting that these methods are limited to linear transformations applied exclusively to the convolutional layers.
Addressing a different yet related problem, (Hager et al., 2023) tackles the scenario where tabular data is assumed to be unavailable during the inference phase but can be utilized during training to leverage semantic dependencies within imaging data. The methodology employs contrastive learning to understand imaging-tabular correspondences. However, as highlighted in Section 1, our work observes that the most relevant tabular features are distinctly present in the imaging data. In certain cases, the tabular data encompasses vital attributes that are challenging or nearly impossible to derive from imaging alone, such as specific protein measurements and metabolic indicators. Moreover, it is worth noting that (Hager et al., 2023) relies on extensive datasets for effective contrastive learning utilization, while our work also addresses situations with smaller datasets lacking comprehensive representation. In such instances, explicitly providing information from the tabular data can prove more beneficial than relying on implicit extraction from the image.
Our work incorporates the concept of conditioning image processing on tabular data by employing a hypernetwork framework. In contrast to other methods, our hypernetwork approach is versatile and comprehensive. Specifically, the proposed hypernetwork learns a general transformation (not necessarily linear) to adapt the parameters transferred to the primary network’s layers, regardless of their architecture or position. This flexibility allows the fusion of imaging with any other modality regardless of the context.
@@ 2.2 Hypernetworks
The notion of hypernetworks, where one network (the hypernetwork) generates the weights and biases for another network (the primary network), was introduced in (Ha et al., 2016). We drew inspiration from two distinct applications of hypernetworks, one being the work of (Littwin and Wolf, 2019). This work utilized 2D images input into a hypernetwork for the reconstruction of depicted 3D objects via the primary network. Another inspiring application is found in the denoising framework presented by (Aharon and Ben-Artzi, 2023), where the hypernetwork was trained to generate primary network parameters conditioned by the latent signal-to-noise ratio (SNR) of the input image.
Hypernetworks have found applications in the medical domain as well. In (Wydmański et al., 2023), a hypernetwork framework processes tabular data using random feature subsets. In this context, a random binary mask selects a subset of tabular features, simultaneously serving as the input to the hypernetwork responsible for generating parameters for the features processing network. However, it is important to note that this approach is unimodal, focusing exclusively on tabular data, and lacks integration with imaging data, a key aspect addressed in our work.
To the best of our knowledge, the utilization of hypernetworks for the fusion of different data modalities in the medical domain remains an unexplored area.
To showcase our approach, we selected two brain MRI analysis challenges that can potentially benefit from the fusion of imaging and tabular data: brain age prediction, a regression problem, and multiclass classification, involving categorization into AD, MCI and CN. Further details about these tasks are provided below.
@@ 2.3 Brain age prediction
The task of predicting human age based on brain anatomy has garnered significant attention in medical research, driven by the development of AI regression tools and the increasing availability of brain imaging data (Franke and Gaser, 2019; Peng et al., 2021; Lee et al., 2022; Cole and Franke, 2017; Feng et al., 2020; Levakov et al., 2020). Current methods, exclusively based on imaging, demonstrate good prediction accuracy, achieving an average gap of less than three years between the actual chronological age and the predicted brain age of healthy subjects. However, since the primary goal of these studies is neural network explainability, aiming to provide insights into brain aging dynamics, relying solely on imaging does not capture other attributes that may influence these processes.
@@
Acknowledging the differences in aging trajectories between males and females (Coffey et al., 1998; Piçarra and Glocker, 2023), we explored whether incorporating subject sex information could enhance brain age prediction. To the best of our knowledge, we are the first to address the task of brain age inference conditioned by the subject’s sex. However, it is essential to note that this exploration is a secondary aim of our study. The primary objective is to showcase and demonstrate our proposed hypernetwork framework, specifically focusing on the fusion of imaging and non-imaging data.
2.4 Alzheimer’s disease classification
@@
Early detection of AD from brain MRI is a topic of active research due to its significant implications. The most challenging aspect lies in classifying individuals with MCI, a transitional stage that serves as a potential predictor for the development of AD. Unlike methods designed for binary classification (CN vs. AD), such as those presented in (Bäckström et al., 2018; Wang et al., 2018), our focus is on multi-class classification into CN, MCI, and AD.
@@
While many studies addressing this problem rely on imaging data alone, as seen in (Wen et al., 2020) and related works, neurology-related literature suggests the inclusion of clinical and demographic data for enhanced prediction. For instance, works like (Letenneur et al., 1999; Fratiglioni et al., 1991) highlight differences in the hazard ratio of AD between women and men, as well as among individuals with varying levels of educational attainment. The Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset offers additional tabular data containing essential attributes and biomarkers for AD detection, such as specific protein measurements in the Cerebrospinal Fluid (CSF) and metabolic indicators derived from Positron Emission Tomography (PET) scans.
@@
These additional attributes and their interactions provide relevant information that may not be present in brain MRI scans alone. In recent years, several multi-modal fusion methods for AD detection have been proposed (Dolci et al., 2022; Venugopalan et al., 2021; El-Sappagh et al., 2020; Zhou et al., 2019; Liu et al., 2018; Spasov et al., 2018). In this manuscript, we specifically refer to (Wolf et al., 2022; Esmaeilzadeh et al., 2018; Prabhu et al., 2022), which addressed the fusion of MRIs and tabular data from the ADNI dataset for three-class AD classification.
@@ 3.1 HyperFusion
Figure 1: A general overview of the proposed hypernetwork framework. The proposed framework ($\mathcal{F}$) consists of a hypernetwork ($\mathcal{H}_{\phi}$) and a primary network ($\mathcal{P}_{\theta}$), where $\phi$ and $\theta$ denote the parameter sets for $\mathcal{H}$ and $\mathcal{P}$, respectively. The primary network parameters $\theta$ are divided into external parameters $\theta_{\mathcal{H}}$ generated by the hypernetwork and internal parameters $\theta_{\mathcal{P}}$ which are learned throughout the backpropagation process. The inputs to $\mathcal{F}$ include the tabular data $T$ which are fed into the hypernetwork and the imaging data $I$ which are fed into the primary network. The entire network compound is trained in an end-to-end manner where the gradients flow through the primary network layers to the hypernetwork layers updating both $\theta_{\mathcal{P}}$ and $\phi$ simultaneously. The external primary network parameters $\theta_{\mathcal{H}}$ are not directly affected by the backpropagation process.
@@
The proposed deep learning framework aims to integrate imaging and tabular data to enhance clinical decision-making. Figure 1 describes its two main building blocks: a hypernetwork denoted by $\mathcal{H}_{\phi}$ and a primary network $\mathcal{P}_{\theta}$, where $\phi$ and $\theta$ represent their respective parameters. For simplicity, we denote the entire network compound as =$\mathcal{F}=\{\mathcal{H}_{\phi},\mathcal{P}_{\theta}\}$.
The input to $\mathcal{F}$ consists of a paired tabular vector of $d$ measurements/values $T\in\mathbb{R}^{d}$ and a multidimensional image $I$, specifically, a 3D MRI. Given a tabular vector $T$, $\mathcal{H}$ produces a data-specific set of network parameters $\theta_{\mathcal{H}}=\mathcal{H}_{\phi}(T)$, which, combined with internally learned parameters $\theta_{\mathcal{P}}$, forms the entire set of primary network parameters $\theta=\{\theta_{\mathcal{H}},\theta_{\mathcal{P}}\}$. The output of the primary network $\mathcal{P}_{\theta}(I)$ represents the desired clinical prediction. During the training phase, the loss backpropagates through both the primary and hypernetwork, influencing both $\theta_{\mathcal{P}}$ and $\phi$. Notably, the external primary network parameters $\theta_{\mathcal{H}}$ are updated indirectly. This structure allows us to condition the image analysis on the tabular data.
@@
Figure 2 provides a detailed view of the proposed architecture and flow. The input image (Figure 2A) is fed into the primary network $\mathcal{P}_{\theta}$ (Figure 2C), a CNN. Simultaneously, the tabular data input (Figure 2A) is processed by the hypernetwork $\mathcal{H}_{\phi}$ (Figure 2B), composed of $K$ sub-networks $\{h_{k}\}_{k=1,\ldots,K}$, each embeds the tabular attributes (see Section 3.1.1) and generates weights and biases $h_{k}(T)=\theta_{h_{k}}$ for a corresponding layer in the primary network $\mathcal{P}_{\theta}$. These external parameters of the primary network’s layers depend on $T$ and influence the extracted imaging features accordingly.
Figure 2: A detailed illustration of the proposed HyperFusion’s network compound $\mathcal{F}$. A: The inputs $T$ and $I$ denote tabular and imaging data, respectively. The images $I$ can be of any type or dimension (in this study we used 3D MRI scans). The tabular data $T\in\mathbb{R}^{d}$ have $d$ attributes. B: The hypernetwork $\mathcal{H}_{\phi}$ is composed of $K$ individual networks, $\{h_{k}\}_{k=1,\ldots,K}$, which generate parameters $h_{k}(T)=\theta_{h_{k}}$ for specific layers of $\mathcal{P}_{\theta}$. Each $h_{k}$ is composed of an MLP embedding network and two linear layers for converting the latent tabular data vector into weights and biases, which are then reshaped to suit their respective mathematical operations in $\mathcal{P}_{\theta}$. C: The primary network $\mathcal{P}_{\theta}$ is a CNN designed for processing the imaging data. The parameters of the primary network $\mathcal{P}_{\theta}$, is composed of $\theta_{\mathcal{H}}=\{\theta_{h_{k}}\}_{k=1,\ldots,K}$ and $\theta_{\mathcal{P}}=\{\theta_{p_{i}}\}_{i\in L}$, where $L$ represents the collection of indexes for layers that are not directly connected to the hypernetwork. The parameters of a specific subset of the primary network’s layers (framed in red) are determined by the hypernetwok. These are the external primary network parameters $\theta_{h_{k}}$. The other primary network parameters $\theta_{p_{i}}$ are internal and are directly updated throughout the backpropagation process.
@@
The dependency on tabular data is determined by the ratio between external and internal parameters. To gain insights, let’s examine two edge cases. When relying solely on internal network parameters (without a hypernetwork), predictions become indifferent of the tabular information. Conversely, the case where all primary network parameters are external and generated by the hypernetwork, approximates a scenario where separate networks are employed for each combination of tabular attributes. For instance, in brain age prediction, the hypernetwork might generate different parameter sets for men and women. However, while there will be some dependencies thanks to the embedding (Section 3.1.1), having all parameters external would probably necessitate an extended training process with a higher amount of training data. In our configurations, we opted for parameters of low-level primary network layers to be internal, assuming that tabular attributes (e.g., sex) are less likely to be relevant for the extraction of the low level imaging features that eventually contribute to the prediction. This chosen combination of parameter sources allows us to maximize the utilization of available training scans while incorporating the corresponding tabular data for additional gains.

### v2

@@ Abstract
The integration of diverse clinical modalities such as medical imaging and the tabular data extracted from patients’ Electronic Health Records (EHRs) is a crucial aspect of modern healthcare. Integrative analysis of multiple sources can provide a comprehensive understanding of the clinical condition of a patient, improving diagnosis and treatment decision. Deep Neural Networks (DNNs) consistently demonstrate outstanding performance in a wide range of multimodal tasks in the medical domain. However, the complex endeavor of effectively merging medical imaging with clinical, demographic and genetic information represented as numerical tabular data remains a highly active and ongoing research pursuit.
We present a novel framework based on hypernetworks to fuse clinical imaging and tabular data by conditioning the image processing on the EHR’s values and measurements. This approach aims to leverage the complementary information present in these modalities to enhance the accuracy of various medical applications. We demonstrate the strength and generality of our method on two different brain Magnetic Resonance Imaging (MRI) analysis tasks, namely, brain age prediction conditioned by subject’s sex and multi-class Alzheimer’s Disease (AD) classification conditioned by tabular data. We show that our framework outperforms both single-modality models and state-of-the-art MRI tabular data fusion methods. A link to our code can be found at https://github.com/daniel4725/HyperFusion.
@@ 1 Introduction
In medical decision-making, clinicians rely on comprehensive patient data, including clinical, genetic, demographic, and imaging information, to gain an in-depth understanding of an individual’s condition, thereby improving diagnostic accuracy and treatment outcomes. Deep Neural Networks (DNNs) consistently achieve state-of-the-art results in a broad spectrum of medical tasks. However, they still fall short of matching human experts’ abilities to integrate information from medical imaging and non-imaging data. Images are high-dimensional, continuous, and spatial, while patients’ Electronic Health Records (EHRs), formatted as numerical, tabular data, typically comprise low-dimensional quantitative attributes with diverse types, scales, and ranges. Vision-Language Models (VLMs), such as CLIP (Contrastive Language–Image Pretraining) [55], have demonstrated remarkable success in capturing cross-modal relationships between textual concepts and images. These methods rely on contrastive learning, where the model is trained to align representations of images with their closely related textual descriptions while distinguishing unrelated pairs. However, in the medical domain, challenges arise due to the non-uniform distribution of data [wang2023sample]. Specifically, there are often no explicit, one-to-one semantic mappings; the same image can be associated with multiple concepts, and a single concept may correspond to different images.
The fusion of imaging and tabular (rather than textual) data presents an additional complication. Although certain measures, such as size or morphology, can be directly correlated with scanned anatomy, many other clinical and demographic attributes, such as medical laboratory results or patient demographics, offer complementary information independent of imaging data. There is no ”correct” or semantic match between a patient’s chest X-ray and their body temperature; the latter, being a continuous variable, can take any value within a given range. However, when combined, both sources of data can be used to improve the assessment of conditions such as pneumonia [Huang2020survey].
@@
To fully exploit both visual and tabular information, we present a conceptually innovative approach that treats clinical measurements and demographic data as priors, influencing the outcomes of an image analysis network. The entire fusion process is analogous to tuning the auditory output of a radio transmitter by manipulating the frequency (selecting stations) and volume. This central concept is implemented through hypernetworks. Originally introduced by [23], the hypernetwork framework consists of a primary network assigned to a specific task and a hypernetwork that generates the weights and biases for specific layers of the primary network. The main strength of this meta-learning framework is that, unlike standard deep models, which remain fixed after training, the primary image-processing network is dynamically adjusted based on the input tabular attributes, even at test time.
The key contributions of this work include: (1) the introduction of HyperFusion, a novel hypernetwork-based framework that effectively fuses imaging and tabular data, where tabular information serves as priors influencing the outcomes of an image analysis network. This dynamic fusion process is shown to be particularly powerful, allowing the image-processing network to be adjusted at test time based on varying input attributes. (2) The application of this framework to two challenging brain MRI analysis tasks: brain age prediction conditioned by sex and classification of Alzheimer’s disease (AD), Mild Cognitive Impairment (MCI), and Cognitively Normal (CN) subjects using their clinical, demographic, and genetic data. (3) Empirical results demonstrating that HyperFusion outperforms state-of-the-art methods in both tasks, providing new insights into the integration of imaging and tabular data in medical diagnostics.
@@
The rest of this paper is organized as follows. Section 2 reviews the fusion of imaging and tabular data, including the concept of hypernetworks and their applications, along with relevant studies on brain age prediction and AD classification. Section 3 introduces our hypernetwork framework, detailing the architectures for each task and key implementation aspects. In Section 4, we describe the experimental setup, including datasets, preprocessing, and training/evaluation procedures, followed by the presentation and discussion of the results. Finally, we conclude in Section 5.
@@ 2.1 Fusion methods
The integration of multiple, diverse medical data modalities is an active area of study [29]. While the fusion of multimodal imaging datasets has been explored for over a decade [45, 7, 63], the merging of imaging and non-imaging data presents additional challenges, and relevant algorithms are relatively recent.
Fusion strategies for combining medical imaging and tabular data can be categorized into three distinct types, as outlined in [Huang2020survey]: early fusion, where original or extracted features are concatenated at the input level, e.g., [9]; joint fusion (or intermediate fusion), where the feature extraction phase is learned as part of the fusion model; and late fusion, where predictions or pre-trained high level features are combined at the decision level, as in [47, 53].
@@
A straightforward and intuitive approach to fuse image and tabular data involves using separate networks for each data type - typically a Multi-Layer Perceptron (MLP) for tabular data and a Convolutional Neural Network (CNN) for imaging. While late fusion methods provide predictions based on collective decisions [53] - joint fusion frameworks embed each modality into a feature vector followed by vector concatenation [16, 14, 65]. However, concatenation methods limit the interaction between tabular data and imaging to high-level descriptors, neglecting fusion in CNN layers where crucial spatial image context is preserved.
More advanced joint fusion techniques apply affine transformation to the image features within the intermediate CNN layers, where the shift and scale parameters are generated by a network that processes another data modality. In [49], a Feature-wise Linear Modulation (FiLM) was introduced to condition image processing on text, applied to a visual question-answering task. Building upon FiLM, [68] introduced the Dynamic Affine Feature Map Transform (DAFT) to fuse brain MRIs and tabular data. However, in these techniques, the interaction between different data types is confined to specific stages of the processing pipeline and linear transformations, potentially limiting the network’s ability to fully exploit the advantages of data fusion.
Addressing a different, yet related problem, [24] tackles the scenario where tabular data is assumed to be unavailable during the inference phase but can be utilized during training to leverage semantic dependencies within imaging data. The methodology employs contrastive learning to understand imaging-tabular correspondences. However, as highlighted in Section 1, our work observes that the most relevant tabular features are distinctly present in the imaging data. In certain cases, the tabular data encompasses vital attributes that are challenging or nearly impossible to derive from imaging alone, such as specific protein measurements and metabolic indicators. Moreover, it is worth noting that [24] relies on extensive datasets for effective contrastive learning utilization, while our work also addresses situations with smaller datasets lacking comprehensive representation. In such instances, explicitly providing information from the tabular data can prove more beneficial than relying on implicit extraction from the image.
Our work incorporates the concept of conditioning image processing on tabular data by employing a hypernetwork framework. In contrast to other methods, our hypernetwork approach is versatile and comprehensive. Specifically, the proposed hypernetwork learns a general transformation (not necessarily linear) to adapt the parameters transferred to the primary network’s layers, regardless of their architecture or position. This flexibility allows for the fusion of imaging with any other modality regardless of the context.
@@ 2.2 Hypernetworks
The notion of hypernetworks, where one network (the hypernetwork) generates the weights and biases for another network (the primary network), was introduced in [23]. We took inspiration from two distinct hypernetwork applications, one of which is the work of [36]. This work utilized 2D images input into a hypernetwork for the reconstruction of depicted 3D objects via the primary network. Another inspiring application is found in the denoising framework presented by [1], where the hypernetwork was trained to generate primary network parameters conditioned by the latent signal-to-noise ratio (SNR) of the input image.
Hypernetworks have also found applications in the medical domain. In [69], a hypernetwork framework processes tabular data using random feature subsets. In this context, a random binary mask selects a subset of tabular features, simultaneously serving as the input to the hypernetwork responsible for generating parameters for the feature processing network. However, it is important to note that this approach is unimodal, focusing exclusively on tabular data, and lacks integration with imaging data, a key aspect addressed in our work.
To the best of our knowledge, the use of hypernetworks for the fusion of different data modalities in the medical domain is proposed here for the first time. To showcase our approach, we selected two brain MRI analysis challenges that can potentially benefit from the fusion of imaging and tabular data: brain age prediction, a regression problem, and multi-class classification, involving categorization into AD, MCI and CN. More details about these tasks are provided below.
@@ 2.3 Brain age prediction
The task of predicting human age based on brain anatomy has garnered significant attention in medical research, driven by the development of AI regression tools and the increasing availability of brain imaging data [18, 48, 33, 11, 17, 35]. Current methods, exclusively based on imaging, demonstrate good prediction accuracy, achieving an average gap of less than three years between the actual chronological age and the predicted brain age of healthy subjects. However, since the primary goal of these studies is the explainability of the neural network, with the aim of providing information on the dynamics of brain aging, relying solely on imaging does not capture other attributes that may influence these processes.
@@
Recognizing the differences in the aging trajectories between men and women [10, 50], we explored whether the incorporation of subject sex information could improve the prediction of brain age. To the best of our knowledge, we are the first to address the task of brain age inference conditioned by the subject’s sex. However, it is essential to note that this exploration is a secondary objective of our study. The primary objective is to showcase and demonstrate our proposed hypernetwork framework, specifically focusing on the fusion of imaging and nonimaging data.
2.4 multi-class AD classification
@@
Early detection of AD from brain MRI is a topic of active research due to its significant implications. The most challenging aspect lies in the classification of individuals with MCI, a transition stage that serves as a potential predictor of the development of AD. Unlike methods designed for binary classification (CN vs AD), such as those presented in [3, 66], our focus is on multiclass classification into CN, MCI and AD.
@@
Although many studies addressing this problem rely only on imaging data, as seen in [67] and related works, the literature related to neurology suggests the inclusion of clinical and demographic data for enhanced prediction. For example, works such as [34, 19] highlight differences in the hazard ratio of AD between women and men, as well as among individuals with varying levels of educational attainment. The Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset offers additional tabular data containing essential attributes and biomarkers for AD detection, such as specific protein measurements in the Cerebrospinal Fluid (CSF) and metabolic indicators derived from Positron Emission Tomography (PET) scans.
@@
These additional attributes and their interactions provide relevant information that may not be present in brain MRI scans alone. In recent years, several multimodal fusion methods have been proposed for AD detection [13, 65, 14, 70, 37, 61]. In this manuscript, we specifically refer to [68, 16, 53], which addressed the fusion of MRIs and tabular data from the ADNI dataset for three-class AD classification.
@@ 3.1 HyperFusion
The proposed deep learning framework aims to integrate imaging and tabular data to enhance clinical decision making. Figure 1 describes its two main building blocks: a hypernetwork denoted by $\mathcal{H}_{\phi}$ and a primary network $\mathcal{P}_{\theta}$, where $\phi$ and $\theta$ represent their respective parameters. For simplicity, we denote the entire network compound by =$\mathcal{F}=\{\mathcal{H}_{\phi},\mathcal{P}_{\theta}\}$.
@@
The input to $\mathcal{F}$ consists of a paired tabular vector of $d$ measurements/values $T\in\mathbb{R}^{d}$ and a multidimensional image $I$, specifically a 3D MRI (Figure 1A). Given a tabular vector $T$, $\mathcal{H}$ produces a data-specific set of network parameters $\theta_{\mathcal{H}}=\mathcal{H}_{\phi}(T)$, which, combined with internally learned parameters $\theta_{\mathcal{P}}$, forms the entire set of primary network parameters $\theta=\{\theta_{\mathcal{H}},\theta_{\mathcal{P}}\}$. The output of the primary network $\mathcal{P}_{\theta}(I)$ represents the desired clinical prediction. In the feedforward pass the input image is fed into the primary network $\mathcal{P}_{\theta}$ (Figure 1C). Simultaneously, the tabular data input is processed by the hypernetwork $\mathcal{H}_{\phi}$ (Figure 1B). The hypernetwork block is composed of $K$ sub-networks $\{h_{k}\}_{k=1,\ldots,K}$, each embeds the tabular attributes (see Section 3.1.1), and generates weights and biases $h_{k}(T)=\theta_{h_{k}}$ for a corresponding layer in the primary network $\mathcal{P}_{\theta}$ (framed in red). These external parameters of the primary network’s layers depend on $T$ and influence the extracted imaging features accordingly.
During the training phase, the loss backpropagates (yellow arrows) through both the primary and hypernetwork, influencing the learnable parameters $\theta_{\mathcal{P}}$ and $\phi$ (marked with yellow stars). Specifically, the gradients update the internal primary network layers and only pass through the external ones to update the parameters in the respective hypernetworks. Notably, the external primary network parameters $\theta_{\mathcal{H}}$ are updated indirectly by the hypernetwork (red arrows). This structure allows us to condition the image analysis on the tabular data.
Figure 1: An illustration of the proposed HyperFusion’s Framework. The two main components - hypernetwork and primary network are shown in the upper and the lower part of the figure, respectively. A: The inputs $T$ and $I$ denote tabular and imaging data, respectively. B: The hypernetwork $\mathcal{H}_{\phi}$ is composed of $K$ individual networks, $\{h_{k}\}_{k=1,\ldots,K}$, which generate parameters $h_{k}(T)=\theta_{h_{k}}$ for specific (external) layers of the primary network $\mathcal{P}_{\theta}$ (red arrows). C: The primary network is composed of internal layers which are updated throughout the backpropagation process (yellow arrows) and external layers (marked in red).
@@
The dependency on tabular data is determined by the ratio between external and internal parameters. To gain insights, let us examine two edge cases. When relying solely on internal network parameters (without a hypernetwork), predictions become indifferent of the tabular information. Conversely, the case where all primary network parameters are external and generated by the hypernetwork, approximates a scenario where separate networks are employed for each combination of tabular attributes. For instance, in brain age prediction, the hypernetwork might generate different parameter sets for men and women. However, while there will be some dependencies thanks to the embedding (Section 3.1.1), having all parameters external would probably necessitate an extended training process with a higher amount of training data. In our configurations, we opted for parameters of low-level primary network layers to be internal, assuming that tabular attributes (e.g., sex) are less likely to be relevant for the extraction of the low level imaging features that eventually contribute to the prediction. This chosen combination of parameter sources allows us to maximize the utilization of available training scans while incorporating the corresponding tabular data for additional gains.